### v1

@@ 2.5 The integral formulas in $q$-spinor variables
where $\gamma_{\mu}$ are the $q$-deformed Dirac matrices defined in (43).
@@
The proof of this result is based on the ideas presented in [8]. Combining (38) with (72) we get
@@
Hence, to solve the integral $\oint_{\Gamma_{q}}D^{q}_{\mu}\Psi$, we will use similary the proof of the Theorem 2.9 of the reference [8], interchanging the Pauli matrices of $q$-deformed Minkowski space (e.g. [14]) by the $q$-deformed Dirac matrices (43), obtaining
@@
We will mention an important consequence of the Definition 3.4. starting of Eqs. (77) and (78): the formulation of the integral $\oint_{\Gamma_{q}}\Psi(u^{\alpha}_{\dot{\beta}}(x))d^{q}x_{\mu}$, which we will mention in the following theorem.
@@
from (38) it follows that
@@
$D^{q}\psi(u^{\alpha}_{\dot{\beta}})-a\phi(u^{\alpha}_{\dot{\beta}})$ = 0absent 0=\ 0= 0 (93)
@@ 3 Differential equations in $q$-spinor variables
We first consider the new $q$-differential operator given by (55). Now, to solve the integral $\oint_{\Gamma_{q}}D^{q}\Psi$, we will use similary the proof of the Theorem 2.9 of the reference [8] (also can see the proof of Proposition 3.10.) to obtain similar expressions to (96) and (97), i.e.
@@ 4 Discussion and suggestions for further work
In Section 2, the equations (51), (60) and (72) describe some $q$-differential operators for $q$-spinor variables. Respect to (51), we can said that the function on the $q$-spinor variables does not depend on ly on the variable $u^{\alpha}_{\dot{\beta}}$ but also on $x_{\nu}$. From this result, the new $q$- differential operator for $q$-spinor variables expressed by (60) was proposed. This operator is motivated from the construction of the any differential operator that satisfy the property $D_{q}^{2}=-\frac{\partial^{2}}{\partial_{q}x^{2}_{\mu}}-\frac{\partial^{2}_{q}}{\partial_{q}x_{\nu}^{2}}$ for all $\mu,\nu=1,2,...,n$. To obtain $D^{2}_{q}$ it is necessary to use the relations (56) and (57). This operator is different to usual Dirac operator and Cauchy Riemann operator mentioned in [3], [4], [6], and [7]. In the case of the Dirac operator for $q$-spinor variables, defined by (72), has been stablished from the Remarks 2.10 and 2.11. This operator is expressed in terms of the $q$-deformed Dirac matrices mentioned by Schmidt [15]. From the $q$-deformed Dirac operator, we define the integral formulas in $q$-spinor variables with the aim to solve the differential equations in $q$-spinor variables. Physically we can said that the Example 3.4 describes the Dirac equation for the electromagnetic case on the $q$-spinor variables, and furthermore the potential $A^{q}_{\mu}(x)$ can be interpreted as the q- electromagnetic potential. There are two further topics arising from this paper which are worth investigation., there is the problem of describing the Maxwell Electrodynamic Algebra which is defined by the following commutation relations

### v2

@@ 2.5 The integral formulas in $q$-spinor variables
where $\gamma_{\mu}$ are the $q$-deformed Dirac matrices defined in the reference [14].
@@
The proof of this result follows from the approach developed in [8], consequently, expression (22) takes the form
@@
Hence, to solve the integral $\oint_{\Gamma_{q}}D^{q}_{\mu}\Psi$, we will use similary the proof of the Theorem 2.9 of the reference [8], interchanging the Pauli matrices of $q$-deformed Minkowski space by the $q$-deformed Dirac matrices (e.g. [14]) , obtaining
@@
We will mention an important consequence of the Proposition 2.15 starting of Eqs. (27) and (28): the formulation of the integral $\oint_{\Gamma_{q}}\Psi(u^{\alpha}_{\dot{\beta}}(x))d^{q}x_{\mu}$, which we will mention in the following theorem.
@@
the term $\frac{\partial^{q}\Psi(u^{\alpha}{\dot{\beta}})}{\partial^{q}u^{\alpha}{\dot{\beta}}}$ may be written in the following form
@@
$D^{q}\psi(u^{\alpha}_{\dot{\beta}})-a\phi(u^{\alpha}_{\dot{\beta}})$ = 0.absent 0=\ 0.= 0 . (43)
@@ 3 Differential equations in $q$-spinor variables
We begin by considering the new $q$-differential operator defined in (5). In order to evaluate the integral $\oint_{\Gamma_{q}}D^{q}\Psi$, we shall proceed in a manner analogous to the proof of Theorem 2.9 in [8] (see also the proof of Proposition 3.10 therein). This approach allows us to derive expressions similar to (46) and (47), namely:
@@ 4 Discussion and suggestions for further work
In Section 2, the equations (1), (10) and (22) describe some $q$-differential operators for $q$-spinor variables. Respect to (1), we can said that the function on the $q$-spinor variables does not depend on ly on the variable $u^{\alpha}_{\dot{\beta}}$ but also on $x_{\nu}$. From this result, the new $q$- differential operator for $q$-spinor variables expressed by (10) was proposed. This operator is motivated from the construction of the any differential operator that satisfy the property $D_{q}^{2}=-\frac{\partial^{2}}{\partial_{q}x^{2}_{\mu}}-\frac{\partial^{2}_{q}}{\partial_{q}x_{\nu}^{2}}$ for all $\mu,\nu=1,2,...,n$. To obtain $D^{2}_{q}$ it is necessary to use the relations (6) and (7). This operator differs from the classical Dirac and Cauchy–Riemann operators discussed in [3, 4, 6, 7, 9], and may be regarded as a $q$-deformed version of the operator $D$ introduced in [9]. In the case of the Dirac operator for $q$-spinor variables, defined by (22), has been stablished from the Remarks 2.10 and 2.11. This operator is expressed in terms of the $q$-deformed Dirac matrices mentioned by Schmidt [15]. From the $q$-deformed Dirac operator, we define the integral formulas in $q$-spinor variables with the aim to solve the differential equations in $q$-spinor variables. Physically we can said that the Example 3.4 describes the Dirac equation for the electromagnetic case on the $q$-spinor variables, and furthermore the potential $A^{q}_{\mu}(x)$ can be interpreted as the q- electromagnetic potential. There are two further topics arising from this paper which are worth investigation., there is the problem of describing the Maxwell Electrodynamic Algebra which is defined by the following commutation relations
@@
The formulation of Dirac operators in $q$-deformed spinorial coordinates constitutes a significant development in the study of quantum theories defined on noncommutative spaces. These structures enable a rigorous extension of the foundations of quantum mechanics and quantum field theory to contexts where classical continuous symmetries are replaced by quantum symmetries, described in terms of Hopf algebras and quantum groups.
Even in its classical form, the Dirac operator plays a central role in the description of spin-$\frac{1}{2}$ particles, such as electrons, and provides a relativistic formulation of quantum mechanics. Its generalisation to the $q$-deformed framework incorporates effects arising from discretisation, quantum curvature, and underlying noncommutative geometries. This generalisation is particularly relevant in theoretical approaches that seek to unify quantum mechanics with gravitational phenomena.
From an algebraic perspective, $q$-deformed Dirac operators are intimately connected with representations of $q$-deformed Clifford algebras and with noncommutative geometries in the sense of Connes. Their study permits the characterisation of quantum manifolds equipped with spinorial structure, thereby paving the way for more general formulations of field theory on curved or quantum spacetimes.
In physical terms, these operators facilitate the modelling of fermionic dynamics in the presence of $q$-deformed external fields, including noncommutative electromagnetic potentials and generalised gauge fields. Notably, the inclusion of a $q$-deformed Maxwell algebra introduces novel forms of interaction between fields and particles, with potential applications to effective models of quantum gravity, quantum cosmology, and condensed matter systems exhibiting discrete or topological symmetries.
The most substantial contribution lies in the capacity of these operators to encapsulate, within a single mathematical framework, the interplay of quantum symmetries, generalised spinorial structures, and noncommutative dynamics. As such, they offer a powerful and versatile tool for probing novel physical regimes in which the geometry of spacetime itself is governed by quantum principles.